OpenStudy (anonymous):

What is the slope between the two points (6, 20) and (2, 4)?

OpenStudy (anonymous):

slope is "change in y" divided by "change in x"

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) when your points are (x1, y1) and (x2, y2)

OpenStudy (anonymous):

y2 is 6 and y1 is 20

OpenStudy (anonymous):

No, choose one point as "point 1" such as point 1 is (6,20), so x1 = 6 and y1 = 20

hero (hero):

Use y = mx + b to solve. Input both points into the formula one at a time to get: 20 = 6m + b 4 = 2m + b

OpenStudy (anonymous):

Or, you can do it that way.

hero (hero):

Next, isolate b in each: b = 20 - 6m b = 4 - 2m

OpenStudy (anonymous):

think so

hero (hero):

Next set b = b 20 - 6m = 4 - 2m

OpenStudy (anonymous):

The possible answers are A. -4 B 4 C 1/4 D 1/4-

hero (hero):

20 - 4 = 6m - 2m 16 = 4m 16/4 = m m = 4

hero (hero):

Sorry for the mistake

OpenStudy (anonymous):

Thanks that's ok.